OpenStudy (anonymous):

Find an equation of the line that satisfies the given condition. (Let x be the independent variable and y be the dependent variable.) The line passing through the origin and parallel to the line passing through the points (2, 5) and (4, 6)

OpenStudy (anonymous):

for da slope i got -1/3?

OpenStudy (campbell_st):

lets check the slope calculation ( 6 - 5)/(4 - 2) = 1/2 does that make sense..

OpenStudy (anonymous):

i dd it 5-6 i got -1 den i did 2- (-1) i got 3

OpenStudy (campbell_st):

(x, y) (4, 6) (2, 5) well the y values are 6 and 5 x values are 4 and 2 is does that make sense...

OpenStudy (anonymous):

lol oh darn i just read wat i ahd put up. sorry its da wrong question!

OpenStudy (campbell_st):

so if you did 5 - 6 = -1 for the y values ( rise) you need to do 2 - 4 = -2 for the x values ( run) slope is rise/run = -1/-2 = 1/2

OpenStudy (campbell_st):

ok...

OpenStudy (anonymous):

yh i got -1/2

OpenStudy (whpalmer4):

Look, plot the two points on your paper. Draw a line through them. Does the line go up, or down? If up, the slope is positive. If down, the slope is negative.

OpenStudy (anonymous):

when i put the answer in, it tells me its wrong @ campbell

OpenStudy (whpalmer4):

Then, write x1 = 4, y1 = 6, x2 = 2, y2 = 5 and plug into the slope formula: \[m=\frac{y_2-y_1}{x_2-x_1}\] \[m=\frac{5-6}{2-4} = \frac{-1}{-2} = \frac{1}{2}\] You have the slope, and you need to force the line to go through a known point, the origin (0,0). \[y-y_0 = m(x-x_0)\]\[y-0 = \frac{1}{2}*(x-0)\]\[y = \frac{1}{2}x\]

OpenStudy (anonymous):

thank u, i forgot to put in da "y" and the "x"